OpenStudy (anonymous):

3^x =2 , 2^y=5 , 5^z= 3 show that xyz=1

OpenStudy (anonymous):

3^x=2 xln3=ln2 x=(ln2)/(ln3) similarly, y=(ln5)/(ln2) z=(ln3)/(ln5) xyz=(ln2)/(ln3) (ln5)/(ln2) (ln3)/(ln5) =1